# msbm10 scaled1095

### v1

@@ 3. Spectra of the compressed shifts on nearly invariant subspaces
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then $A_{v}$ is unitarily equivalent to the compressed shift $S_{\theta_{v}}$ on a model space $K_{\theta_{v}}$, where $\theta_{v}$ is a Frostman shift of $\theta$.
It is shown in [8] (see also [6, Proposition 4.3] or [21, Theorem 13.1]) that for an inner function $\theta$ and $v\in\mathbb{D}$, the Crofoot transform
@@
By the unitary mapping $J_{v}:\;K_{\theta}\rightarrow K_{\theta_{v}},$ the following commutative diagram is derived in [21, Lemma 13.3]: that is, $J_{v}A_{v}=S_{\theta_{v}}J_{v}$.
@@
For the compressed shift $S_{\theta}$ on $K_{\theta}$, [21, Lemma 2.5] or [11, Corollary 9.24] exhibits its point spectrum as
@@
Meanwhile, the corresponding eigenvectors are found in [11, Corollary 9.25] by showing
$\ker(S_{\theta}-\lambda I)=\mathbb{C}Ck_{\lambda}^{\theta},\;\ker(S_{\theta}^{*}-\overline{\lambda}I)=\mathbb{C}k_{\lambda}^{\theta}$ (3.7)
with conjugation $C$ in (1.1) and $\lambda\in\mathbb{D}$ satisfying $\theta(\lambda)=0.$ So we can use them to describe the point spectrum and corresponding eigenvectors of $A_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$ by (3.5).
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then the point spectrum is given by
@@
where $S_{\theta_{v}}$ denotes the compressed shift on $K_{\theta_{v}}.$ This immediately yields $\sigma_{p}(A_{v})=\sigma_{p}(S_{\theta_{v}}).$
Applying (3.6) to the Frostman shift $\theta_{v}$ defined in (3.3), we obtain
@@
This has also been established in [21, Theorem 14.1].
@@
Applying $B_{v}=(A_{v})^{*}$ and $B_{v}=CA_{v}C$, there follows a remark on $B_{v}$’s point spectrum.
@@
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then the point spectrum is given by
@@
Two illustrative examples demonstrate the eigenvalues and eigenvectors of $A_{v}$ and $B_{v}$.
@@
(1)  Let $\theta(z)=z^{2}\frac{z-a}{1-\overline{a}z}$ when $10<|a|<$ 1 . Since $K_{\theta}=K_{z}\oplus zK_{z\frac{z-a}{1-\overline{a}z}},$ we take the basis $e_{1}(z)=1,\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z^{2}/(1-\overline{a}z).$ Then it holds that
@@
Its eigenvalues satisfy $(\lambda-a)\lambda^{2}-v(1-\lambda\overline{a})=0$ or $\theta(\lambda)=v.$ And let $[B_{v}]=[A_{v}]^{*}$, we deduce that
$([B_{v}]-\overline{\lambda}I)k_{\lambda}^{\theta}$ $=\left(\begin{array}[]{ccc}-\overline{\lambda}&1&-{a}\\ 0&-\overline{\lambda}&1\\ \overline{v}&0&-\overline{\lambda}+\overline{a}\end{array}\right)\left(\begin{array}[]{c}1\\ \frac{\overline{\lambda}}{1-a\overline{\lambda}}\\ \frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\end{array}\right)=\left(\begin{array}[]{c}-\overline{\lambda}+\frac{\overline{\lambda}}{1-a\overline{\lambda}}-\frac{a\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ -\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}+\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ \overline{v}+\frac{\overline{\lambda}^{2}(\overline{a}-\overline{\lambda})}{1-a\overline{\lambda}}\end{array}\right)=0$
@@
(2)  Let $\theta(z)=z\left(\frac{z-a}{1-\overline{a}z}\right)^{2}$ when $10<|a|<$ 1 . Since $K_{\theta}=K_{z\frac{z-a}{1-\overline{a}z}}\oplus z\frac{z-a}{1-\overline{a}z}K_{\frac{z-a}{1-\overline{a}z}},$ the basis is $e_{1}(z)=1/(1-\overline{a}z),\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z(z-a)/(1-\overline{a}z)^{2}.$ Then we deduce that
@@
Meanwhile, for the adjoint $[B_{v}]=[A_{v}]^{*}$, it holds that
@@
$=\left(\begin{array}[]{c}0\vspace{1mm}\\ -a\overline{v}+\overline{\lambda}\frac{\overline{a}-\overline{\lambda}}{1-a\overline{\lambda}}\frac{1-a\overline{\lambda}-1+|a|^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ \overline{v}-\overline{\lambda}\frac{(\overline{\lambda}-\overline{a})^{2}}{(1-a\overline{\lambda})^{2}}\end{array}\right)=\left(\begin{array}[]{c}0\vspace{1mm}\\ -a\overline{v}+a\overline{\lambda}\frac{(\overline{a}-\overline{\lambda})^{2}}{(1-a\overline{\lambda})^{2}}\vspace{1mm}\\ 0\end{array}\right)=0$
@@
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then $\sigma_{p}(A_{z}^{\mathcal{M}})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}$ with $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. For each $\lambda\in\sigma_{p}(A_{z}^{\mathcal{M}}),$ the corresponding eigenvector is $h(Ck_{\lambda}^{\theta})=h\widetilde{k_{\lambda}^{\theta}}.$
@@
which is a special case of [21, Theorem 14.1] with the inner function vanishing at zero. Therefore, the preceding analysis enables us to derive a complete description for the spectrum of $A_{z}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
where $v=\langle\theta,|h|^{2}\rangle$ with $|v|<1$.
@@ 4. Invariant subspaces of the compressed shifts on nearly invariant subspaces
In this section, we aim to characterize the invariant subspaces of our c.n.u. operator $A_{v}$ or $B_{v}$ with $|v|<1$ on a model space $K_{\theta}$ with $\theta(0)=0$. As an application, we completely solve Question 1 by describing all the invariant subspaces of the compressed shift $A_{z}^{\mathcal{M}}$ on arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
Theorem A [9, Theorem 3.1] For inner $\phi$ and $\theta$ and $g\in H^{2},$ the following are equivalent:
@@
Let $\mathcal{N}$ be an invariant subspace of $B_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$, then $\mathcal{N}$ is nearly $S^{*}$-invariant and then $\mathcal{N}=gK_{\phi}$ is an invariant subspace for $B_{v}$ if and only if $\theta\overline{g\phi}\in H^{2}$ and $(B_{v}g)/g\in K_{\phi}$.
@@
Let $\theta(z)=z\frac{z-a}{1-\overline{a}z}$ when $10<|a|<$ 1 . Write $e_{1}(z)=1/(1-\overline{a}z)$ and $e_{2}(z)=z/(1-\overline{a}z).$ So $A_{v}e_{1}=e_{2}$ and $A_{v}e_{2}=ve_{1}+(a-v\overline{a})e_{2}.$ Therefore $A_{v}$ behaves as the matrix
@@
$(B_{v}-\overline{\lambda}I)^{n+1}(k_{\lambda}^{\theta})^{(n)}(z)$ $=\sum_{i=0}^{n+1}C_{n+1}^{i}(-\overline{\lambda})^{n+1-i}B_{v}^{i}(k_{\lambda}^{\theta})^{(n)}(z)$
@@
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Suppose the eigenvalue $\overline{\lambda}$ of $B_{v}$ has multiplicity $n+1$ for $n\geq 2$ and satisfies $\theta(\lambda)=v,$ then $(k_{\lambda}^{\theta})^{(n)}=\frac{n!z^{n}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}}$ is a generalized eigenvector of $B_{v}$.
@@
As an example, consider the singular inner function $u(z)=\exp(\frac{z-1}{z+1})$, which satisfies $u(0)=1/e$ and never vanishes. Defining the Frostman shift of the inner function $u$ as
@@
we observe $\theta$ is inner, $\theta(0)=0$, and $\theta$ never takes the value $v:=-1/e$. If the spectrum of $A_{v}$ or $B_{v}$ is disconnected then it always has invariant subspaces. But the spectrum in this example is one point on the circle, so how do we answer Question 2?
At this moment, the unitary equivalence in the diagram (3.5), together with, the fact the invariant subspaces of $S_{\theta_{v}}$ in $K_{\theta_{v}}$ are precisely $K_{\theta_{v}}\ominus K_{\phi}$, where $\phi$ divides $\theta_{v}$ (see, e.g.[11, Proposition 9.14]) establish the invariant subspaces for $A_{v}$ as below.
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$, then the invariant subspaces of $A_{v}$ in $K_{\theta}$ are precisely $J_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $\phi$ divides $\theta_{v}$ and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is defined in (3.4).
@@
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then every $A_{z}^{\mathcal{M}}$-invariant subspace of $\mathcal{M}$ is of the form $hJ_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $v=\langle\theta,|h|^{2}\rangle$ with $|v|<1$, $\phi$ is an inner function dividing the Frostman shift $\theta_{v}$ of $\theta$, and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is the inverse of the Crofoot transform.
To conclude this paper, we present a concrete example to illustrate our main results.
@@
This yields that $p=2/3$ and $q=-1/3$. Now $\|k\|^{2}=\langle k,k\rangle=k(0)=2/3$, leading to the normalized reproducing kernel
@@
Now the model space $K_{\theta}$ must contain $1/(2-z)$ and $z/(2-z)$, which enables us to choose
@@
So $(1+z)K_{z^{2}}=hK_{\theta}$ with extremal function $h$ in (4.2) and associated inner function $\theta$ in (4.3).
We then compute
@@
and Theorem 3.8 ensures the functions $(z-a)(z+1)$ and $(z-b)(z+1)$ are the eigenvectors corresponding to the eigenvalues $b$ and $a$, respectively.

### v2

@@ 3. Spectra of the compressed shifts on nearly invariant subspaces
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then $A_{v}$ is unitarily equivalent to the compressed shift $S_{\theta_{v}}$ on a model space $K_{\theta_{v}}$, where $\theta_{v}$ is a Frostman shift of $\theta$.
It is shown in [9] (see also [7, Proposition 4.3] or [24, Theorem 13.1]) that for an inner function $\theta$ and $v\in\mathbb{D}$, the Crofoot transform
@@
By the unitary mapping $J_{v}:\;K_{\theta}\rightarrow K_{\theta_{v}},$ the following commutative diagram is derived in [24, Lemma 13.3]: that is, $J_{v}A_{v}=S_{\theta_{v}}J_{v}$.
@@
For the compressed shift $S_{\theta}$ on $K_{\theta}$, [24, Lemma 2.5] or [12, Corollary 9.24] exhibits its point spectrum as
@@
Meanwhile, the corresponding eigenvectors are found in [12, Corollary 9.25] by showing
$\ker(S_{\theta}-\lambda I)=\mathbb{C}Ck_{\lambda}^{\theta}\;\mbox{and}\;\ker(S_{\theta}^{*}-\overline{\lambda}I)=\mathbb{C}k_{\lambda}^{\theta}$ (3.7)
with conjugation $C$ in (1.1) and $\lambda\in\mathbb{D}$ satisfying $\theta(\lambda)=0.$ So we further proceed to describe the point spectrum and corresponding eigenvectors of $A_{v}$ on $K_{\theta}$ with $\theta(0)=0$ by (3.5).
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then its point spectrum is given by
@@
where $S_{\theta_{v}}$ denotes the compressed shift on $K_{\theta_{v}}.$ This immediately yields $\sigma_{p}(A_{v})=\sigma_{p}(S_{\theta_{v}}).$ Applying (3.6) to the Frostman shift $\theta_{v}$ defined in (3.3), we deduce
@@
This has also been established in [24, Theorem 14.1].
@@
Applying $B_{v}=(A_{v})^{*}$ and $B_{v}=CA_{v}C$, a remark follows on $B_{v}$’s point spectrum.
@@
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then the point spectrum is given by
@@
Two illustrative examples are provided to demonstrate the eigenvalues and eigenvectors of $A_{v}$ and $B_{v}$.
@@
(1)  Let $\theta(z)=z^{2}\frac{z-a}{1-\overline{a}z}$ when $0<|a|<1.$ Since $K_{\theta}=K_{z}\oplus zK_{z\frac{z-a}{1-\overline{a}z}},$ we take the basis $e_{1}(z)=1,\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z^{2}/(1-\overline{a}z).$ Then it holds that
@@
Its eigenvalues satisfy $(\lambda-a)\lambda^{2}-v(1-\lambda\overline{a})=0$ or $\theta(\lambda)=v.$ And let $B_{v}]=[A_{v}]^{*}$, we deduce that
$([B_{v}]-\overline{\lambda}I)k_{\lambda}^{\theta}$ $=\left(\begin{array}[]{ccc}-\overline{\lambda}&1&-{a}\\ 0&-\overline{\lambda}&1\\ \overline{v}&0&-\overline{\lambda}+\overline{a}\end{array}\right)\left(\begin{array}[]{c}1\\ \frac{\overline{\lambda}}{1-a\overline{\lambda}}\\ \frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\end{array}\right)=\left(\begin{array}[]{c}-\overline{\lambda}+\frac{\overline{\lambda}}{1-a\overline{\lambda}}-\frac{a\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vskip 2.84526pt\\ -\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}+\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vskip 2.84526pt\\ \overline{v}+\frac{\overline{\lambda}^{2}(\overline{a}-\overline{\lambda})}{1-a\overline{\lambda}}\end{array}\right)=0$
@@
(2)  Let $\theta(z)=z\left(\frac{z-a}{1-\overline{a}z}\right)^{2}$ when $0<|a|<1.$ Since $K_{\theta}=K_{z\frac{z-a}{1-\overline{a}z}}\oplus z\frac{z-a}{1-\overline{a}z}K_{\frac{z-a}{1-\overline{a}z}},$ the basis is $e_{1}(z)=1/(1-\overline{a}z),\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z(z-a)/(1-\overline{a}z)^{2}.$ Then we deduce that
@@
Meanwhile, for the adjoint $B_{v}]=[A_{v}]^{*}$, it holds that
@@
$=\left(\begin{array}[]{c}0\vskip 2.84526pt\\ -a\overline{v}+\overline{\lambda}\frac{\overline{a}-\overline{\lambda}}{1-a\overline{\lambda}}\frac{1-a\overline{\lambda}-1+|a|^{2}}{1-a\overline{\lambda}}\vskip 2.84526pt\\ \overline{v}-\overline{\lambda}\frac{(\overline{\lambda}-\overline{a})^{2}}{(1-a\overline{\lambda})^{2}}\end{array}\right)=\left(\begin{array}[]{c}0\vskip 2.84526pt\\ -a\overline{v}+a\overline{\lambda}\frac{(\overline{a}-\overline{\lambda})^{2}}{(1-a\overline{\lambda})^{2}}\vskip 2.84526pt\\ 0\end{array}\right)=0$
@@
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then $\sigma_{p}(A_{z}^{\mathcal{M}})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}$ with $v=\langle\theta,|h|^{2}\rangle$. For each $\lambda\in\sigma_{p}(A_{z}^{\mathcal{M}}),$ the corresponding eigenvector is $h(Ck_{\lambda}^{\theta})=h\widetilde{k_{\lambda}^{\theta}}$.
@@
which is a special case of [24, Theorem 14.1] with the inner function vanishing at zero. Therefore, the preceding analysis enables us to derive a complete description for the spectrum of $A_{z}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
where $v=\langle\theta,|h|^{2}\rangle$.
@@ 4. Invariant subspaces of the compressed shifts on nearly invariant subspaces
In this section, we aim to characterize the invariant subspaces of our c.n.u. operator $A_{v}$ or $B_{v}$ on a model space $K_{\theta}$ with $\theta(0)=0$. As an application, we completely solve Question 1 by describing all the invariant subspaces of the compressed shift $A_{z}^{\mathcal{M}}$ on arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
@@
Theorem A [10, Theorem 3.1] For inner $\phi$ and $\theta$ and $g\in H^{2},$ the following are equivalent:
@@
Let $\mathcal{N}$ be an invariant subspace of $B_{v}$ on $K_{\theta}$ with $\theta(0)=0$, then $\mathcal{N}$ is nearly $S^{*}$-invariant and then $\mathcal{N}=gK_{\phi}$ is an invariant subspace for $B_{v}$ if and only if $\theta\overline{g\phi}\in H^{2}$ and $(B_{v}g)/g\in K_{\phi}$.
@@
Let $\theta(z)=z\frac{z-a}{1-\overline{a}z}$ when $0<|a|<1.$ Write $e_{1}(z)=1/(1-\overline{a}z)$ and $e_{2}(z)=z/(1-\overline{a}z).$ So $A_{v}e_{1}=e_{2}$ and $A_{v}e_{2}=ve_{1}+(a-v\overline{a})e_{2}.$ Therefore $A_{v}$ behaves as the matrix
@@
$\quad\quad(B_{v}-\overline{\lambda}I)^{n+1}(k_{\lambda}^{\theta})^{(n)}(z)$
$=\sum_{i=0}^{n+1}C_{n+1}^{i}(-\overline{\lambda})^{n+1-i}B_{v}^{i}(k_{\lambda}^{\theta})^{(n)}(z)$
@@
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Suppose the eigenvalue $\overline{\lambda}$ of $B_{v}$ has multiplicity $n+1$ for $n\geq 2$ and satisfies $\theta(\lambda)=v,$ then $(k_{\lambda}^{\theta})^{(n)}=\frac{n!z^{n}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}}$ is a generalized eigenvector of $B_{v}$.
@@
As an example, consider the singular inner function $u(z)=\exp(\frac{z-1}{z+1})$, which satisfies $u(0)=1/e$ and never vanishes. Defining its Frostman shift of $u$ as
@@
we observe $\theta$ is inner with $\theta(0)=0$ that never attains the value v:=1/ev:=-1/e. If the spectrum of $A_{v}$ or $B_{v}$ is disconnected, invariant subspaces necessarily exist. In this example, however, the spectrum consists of a single point on the circle, leaving Question 2 unsolved.
At this stage, the unitary equivalence illustrated in the diagram (3.5) becomes crucial. Together with the fact that the invariant subspaces of $S_{\theta_{v}}$ in $K_{\theta_{v}}$ are precisely those of the form $K_{\theta_{v}}\ominus K_{\phi}$ for inner divisors $\phi$ of $\theta_{v}$ (see, e.g.[12, Proposition 9.14]). This equivalence yields an explicit description of the invariant subspaces for $A_{v}$, which we now present.
@@
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$, then the invariant subspaces of $A_{v}$ in $K_{\theta}$ are precisely $J_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $\phi$ divides $\theta_{v}$ and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is defined in (3.4).
@@
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then every $A_{z}^{\mathcal{M}}$-invariant subspace of $\mathcal{M}$ is of the form $hJ_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $v=\langle\theta,|h|^{2}\rangle$, $\phi$ is an inner function dividing the Frostman shift $\theta_{v}$ of $\theta$, and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is the inverse of the Crofoot transform.
Concluding this paper, we present an example to illustrate our main results.
@@
This yields that $p=2/3$ and q=1/3q=-1/3. Now $\|k\|^{2}=\langle k,k\rangle=k(0)=2/3$, leading to the normalized reproducing kernel
@@
Now the model space $K_{\theta}$ must contain 1/(2z)1/(2-z) and z/(2z)z/(2-z), which enables us to choose
@@
So $(1+z)K_{z^{2}}=hK_{\theta}$ with extremal function $h$ in (4.2) and associated inner function $\theta$ in (4.3). We then compute
@@
and Theorem 3.8 ensures the functions (za)(z+1)(z-a)(z+1) and (zb)(z+1)(z-b)(z+1) are the eigenvectors corresponding to the eigenvalues $b$ and $a$, respectively.
@@
Data Availability   Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest   The authors state that there is no conflict of interest.